Two concentric spheres 210 mm and 300 mm diameters with the space between

them evacuated are to be used to store liquid air (– 153°C) in a room at 27°C. The surfaces of the

spheres are flushed with aluminium (ε = 0.03) and latent heat of vaporization of liquid air is

209.35 kJ/kg. Calculate the rate of evaporation of liquid air.


1
Expert's answer
2021-09-24T09:25:34-0400

"Q=\frac{A_1\sigma(T_1^4-T_2^4)}{\frac{1-\varepsilon_1}{\varepsilon_1}+1+(\frac{1-\varepsilon_1}{\varepsilon_1})\frac{A_1}{A_2}}=1.26~W,"

"v=\frac Qh=\frac{1.26\cdot 3600}{209.35\cdot 1000}=21.7~\frac g{hour}."